# Spiking mode-based neural networks

### v2

@@ Abstract
Spiking neural networks play an important role in brain-like neuromorphic computations and in studying working mechanisms of neural circuits. One drawback of training a large scale spiking neural network is that updating all weights is quite expensive. Furthermore, after training, all information related to the computational task is hidden into the weight matrix, prohibiting us from a transparent understanding of circuit mechanisms. Therefore, in this work, we address these challenges by proposing a spiking mode-based training protocol, where the recurrent weight matrix is explained as a Hopfield-like multiplication of three matrices: input, output modes and a score matrix. The first advantage is that the weight is interpreted by input and output modes and their associated scores characterizing the importance of each decomposition term. The number of modes is thus adjustable, allowing more degrees of freedom for modeling the experimental data. This significantly reduces the training cost because of significantly reduced space complexity for learning. Training spiking networks is thus carried out in the mode-score space. The second advantage is that one can project the high dimensional neural activity (filtered spike train) in the state space onto the mode space which is typically of a low dimension, e.g., a few modes are sufficient to capture the shape of the underlying neural manifolds. We successfully apply our framework in two computational tasks—digit classification and selective sensory integration tasks. Our method accelerate the training of spiking neural networks by a Hopfield-like decomposition, and moreover this training leads to low-dimensional attractor structures of high-dimensional neural dynamics.
@@ I Introduction
We review relevant existing methods of training spiking neural networks and highlight the difference from ours. Recurrent neural networks (RNNs) using continuous rate dynamics of neurons can be used to generate coherent output sequences [7], in which the network weights are trained by the first-order reduced and controlled error (FORCE) method, where the inverse of the rate correlation matrix is iteratively estimated to update the weights during learning [7, 5]. This method can be generalized to supervised learning in spiking neural networks [8, 9, 10]. A recent work proposed to train the rate network first and then rescale the synaptic weights to adapt to a spiking setting [11], while earlier works proposed to map a trained continuous-variable rate RNN to a spiking RNN model [12, 13]. These methods rely either on standard efficient methods like FORCE for rate models, or on heuristic strategies to modify the weights on the rate counterpart. Another recently proposed route is a non-linear voltage-based three-factor learning rule for multilayer networks of deterministic integrate-and-fire neurons [14]. All these studies involve in real valued weights. In contrast, our current SMNN framework works directly on the mode space of the recurrent connections underlying the spiking dynamics for which many biological plausible factors can be incorporated, e.g., membrane and synapse time scales, cell types and refractory time etc.
@@
In essence, we decompose the traditional real-valued weights as three matrices: the left one acts as the input mode space, the right one acts as the output mode space, and the middle one encodes the importance of each mode in the corresponding space. Therefore, we can interpret the weight as two mode spaces and a score matrix. The neural dynamics can thus be projected to the input mode space which is an intrinsically low dimensional space. In addition, the leaky integrated fire (LIF) model can be discretized into a discrete dynamics with intrinsic synaptic dynamics and threshold-type firing, for which a surrogate gradient descent with an additional steepness hyperparameter can be applied [6, 27]. We then show in experiments how the SMNN framework is successfully applied to classify pixel-by-pixel MNIST datasets and furthermore to context dependent computation in neuroscience. For both computational tasks, we reveal attractor structures in a low-dimensional mode space (in the number of modes) onto which the high dimensional neural dynamics can be projected. Overall, we construct a computational efficient (less model parameters required because of a few dominating modes) and conceptually simple framework to understand challenging spike-based computations (attractor picture in physics). This finding will inspire more physics researches on the high dimensional spiking neural dynamics including learning.
@@ II.1 Recurrent spiking dynamics with mode-based connections
Before specifying the neural dynamics, we remark a few differences from the singular value decomposition method. First, the mode decomposition learning rule is inspired in physics by Hopfield model (a toy model of associative memory [30], see also our two recent papers [24, 26]). In the generalized form [26], one can flexibly interpret the rank number as the number of patterns, and this number is adjustable. Second, the elements of the importance matrix is not necessarily positive, and can be learned during the training protocol. Third, the input and output mode vectors are not necessarily orthogonal, allowing for more degrees of freedom to interpret experimental data or execute computational tasks. In addition, these mode vectors help to project the dynamics trajectories in the high-dimensional state space onto the low-dimensional mode space (the dimensionality is determined by $P\ll N$).
@@
where $\tau_{r}$ and $\tau_{d}$ refer to the synaptic rise time and the synaptic decay time, respectively; $t_{i}^{k}$ refers to the $k$-th spiking time of unit $i$. Different time scales of synaptic filters are due to different types of receptors (such as fast AMPA, relatively fast GABA, and slow NMDA receptors) with different temporal characteristics [1]. Note that the above synaptic filter has also been used in a previous work [11], which proposed to train the rate network first and then rescale the synaptic weights to adapt to a spiking setting. We remark that other types of synaptic filter can also be used, e.g. the simple exponential type used in previous works [9, 27].
@@ II.2.1 MNIST data learning
The MNIST dataset consists of handwritten digits, each of which is composed of 28x28 pixels, belonging to $10$ different classes (6000 images for each class). The dataset is commonly used as a benchmark classification for neural networks [32]. The handwritten digit is a static image, and thus must be transformed into spiking time series. We thus introduce an additional transformation layer consisting of spiking neurons before the hidden layer. More precisely, each image is converted into spiking activity using one neuron per pixel. The transformation layer is then modeled by a population of Poisson spiking neurons with a maximum firing rate $f_{\rm max}$. Each neuron in this layer generates a Poisson spike train with a rate $f_{i}=\frac{g_{i}}{255}f_{\rm max}$, where $g_{i}$ is the corresponding pixel intensity. We present the same static image every $0.2\,{\rm{ms}}$ (step size) and for a total of $100$ time steps as input to the transformation layer (therefore for a duration $20\ {\rm{ms}}$), to convert the image into spiking activity input to the hidden layer [Figure 1 (a) and (b)].
@@ III Simulation results of mode decomposition learning
where a steepness parameter $k$ = 25. This is because, $\lim_{k\to\infty}\Theta^{\prime}(x)=\delta(x)$, where $x=U-U_{\rm thr}$. In the forward pass, the original form of the step function is used. Other parameters include $\Delta t=0.2\,{\rm{ms}}$, $\tau_{\rm mem}=20\,{\rm{ms}}$, $\tau_{d}=30\,{\rm{ms}}$, $\tau_{r}=2\,{\rm{ms}}$, and the refractory period $t_{\rm ref}=2\,{\rm{ms}}$ for all tasks. These hyper-parameters are summarized in Table 3 (see Appendix C). Because we have not imposed sparsity constraint on the connectivity, we take the initialization scheme that $[\bm{\xi}^{\rm in}\bm{\Sigma}(\bm{\xi}^{\rm out})^{\top}]_{ij}\sim O(\frac{1}{\sqrt{N}})$, similar to what is done in multi-layered perceptron learning [26]. The training is implemented by the adaptive-moment-estimation (Adam [34]) based stochastic gradient descent algorithm that minimizes the loss function (cross-entropy for classification or mean-squared error for regression in the second task) with a learning rate of $0.001$. We remark that training in the mode space can be done by using automated differentiation on PyTorch. However, we still leave the detailed derivation of the learning rule to Appendix B. Codes are available in our GitHub [35].
@@ III.1 MNIST classification task
Figure 2: Learning performance of MNIST classification task. Five independent runs are used to estimate the standard deviation. (a) Test accuracy versus different mode size. Spiking model without mode decomposition learning (this counterpart is called SNN) is compared. SMNN indicates the mode-based learning of spiking networks. (b) Loss function as a function of training mini-batch, each composed of $100$ digit images. The mode size varies, and the network size $N=100$. (c) Comparison of the accuracies between rate and spiking models with the same fixed network size $N$=200. Rate networks with mode-decomposition learning (MDL RNN) and without MDL (RNN) are also considered. (d) Membrane potential trace for three typical reservoir neurons in response to spike train inputs after training. The triangle or square marks when the filtered spike train takes a maximum value. $(P,N)=(3,100)$. Note that three moments for the maximal firing rates overlap with each other, but this is not the case for other neurons. (e) Spike trains of reservoir neurons with neuron firing rate (right) and population firing fraction (bottom) in response to an input image of digit $2$. $(P,N)=(3,100)$.
@@ III.2 Contextual integration task
As shown in Figure 3 (a), the test MSE [i.e., $\mathcal{L}^{2}$, see Eq. (11)] for the SMNN decreases with the mode size. In particular, a larger neural population is better. Even for a few modes like $P=3$, the predicted output already matches the target very well [see Figure 3 (c)]. The SNN without the mode decomposition achieves a lower MSE, but we emphasize that the accuracies behind the MSE within the shown range (from $0.01$ to $0.02$) do not look very different [see Figure 3 (c) for the case of $P=3$]. In this sense, compared to SNN, SMNN saves the training parameters. The training dynamics in Figure 3 (b) shows that a larger value of $P$ speeds up the learning. Figure 3 (c) confirms that our training protocol succeeds in reproducing the result of Monkey’s experiments on the task of flexible selective integration of sensory inputs. We can even look at the dynamics profile of the membrane potential [Figure 3 (d)]. During the stimulus period, each neuron has its own time scale to encode the input signals, and at a network scale, we do not observe any regular patterns but an asynchronous pattern typically observed in prefrontal cortex. After the sensory input is turned off, the network is immediately required to make a decision, and we observe that the firing frequency of the neural pool is elevated already a bit earlier than the moment of removing input, but a very low population firing rate is still maintained [Figure 3 (d)]. This observation is a reflection of neural dynamics on the neural manifolds underlying the perceptual decision making, which we shall detail below.
@@ III.3 Power law for connectivity importance scores
Figure 4: Connectivity importance $|\lambda_{\mu}|$ versus rank (in descending order). Both MNIST classification and contextual integration task are considered, and the simulation conditions are the same with that in Figure 2 for MNIST and that in Figure 3 for context dependent computation. We fix $P=30$. Three independent runs are used to estimate the standard deviation. We also define a more precise measure $\tau_{\mu}=\chi\|\bm{\xi}^{\rm in}_{\mu}\|_{2}+\chi\|\bm{\xi}^{\rm out}_{\mu}\|_{2}+|\lambda_{\mu}|$, where $\chi=\sum_{\mu}|\lambda_{\mu}|/\sum_{\mu}(\|\bm{\xi}^{\rm in}_{\mu}\|_{2}+\|\bm{\xi}^{\rm out}_{\mu}\|_{2})$. There appears the piecewise power law behavior in the inset of the log-log plot. The colored vertical dashed lines mark the fitting ranges for different network sizes. (a) MNIST classification. (b) Contextual integration task.
@@ III.4 Reduced dynamics in the mode space
Here, we will show how the dynamics of network activity based on our SMNN learning can be projected to a low-dimensional counterpart. We follow the previous works of low-rank recurrent neural networks [21, 22]. To construct a subspace spanned by orthogonal bases, we first split $\bm{W}^{\rm in}$ into the parts parallel and orthogonal to the input mode $\bm{\xi}^{\rm in}$ (or the output mode),
@@ III.6 Spiking variability measured by Fano factor
The spiking activity is commonly irregular, displaying strong variability even when the neural population is shown identical stimuli. One can thus count the spiking events in a time window of duration $T$ for one neuron, and then repeat the experiment over many trials and finally measure the mean and variance of the counts. The Fano factor is then calculated as [1]
@@
We plot the Fano factor distribution across neurons in Fig. 7. The Fano factor is estimated over the entire experimental window. For the MNIST classification task, the Fano factor is mainly distributed from $2$ to $4$ [Figure 7 (a)], indicating highly unreliable firing patterns, which corresponds to that observed in Figure 2 (e). For an input image, some neurons fire at high frequencies, while others remain silent. This non-Poissonian phenomenon has also been observed in cortical areas [36]. For the context-integration task, the Fano factor is surprisingly concentrated around one [Figure 7 (b)], except for a long tail at large values of Fano factor. This model observation highlights the potential of our methods in studying the alignment between the simulation and neural experiments of context dependent cognitive decision making.
@@ III.7 Ablation studies on the refractory, activation function and network robustness
where $\beta=20$ to mimic step function. Note that in this control experiment, we use this transfer function both in forward and backward passes during the training. As shown in Figure 8 (a) for the MNIST task, the absence of the refractory period allows the network to respond more correctly to external stimuli. Performance without the refractory period is slightly better than with the refractory period, but as the mode number increases, the gap shrinks. For the contextual integration task, an appropriate refractory period is able to optimize network performance [Figure 8 (b)], consistent with previous results in standard SNNs [11]. Finally, we remark that networks that use a smooth transfer function instead of the step function perform much worse than spiking networks using a sharp step function (neurons can thus fire or be silent), indicating the importance of spiking event.
@@
We next consider the network robustness in a weight-pruning experiment. More precisely, for a well-trained connectivity matrix $W^{\rm rec}_{ij}=\sum_{\mu}\lambda_{\mu}\xi^{\rm in}_{i\mu}\xi^{\rm out}_{j\mu}$, we set some elements where their absolute values $|W^{\rm rec}_{ij}|<\theta$ to zero, with $\theta$ is an adjustable pruning threshold. This operation leads to a diluted connectivity matrix $\hat{\bm{W}}^{\rm rec}$:

### v3

@@ Abstract
Spiking neural networks play an important role in brain-like neuromorphic computations and in studying working mechanisms of neural circuits. One drawback of training a large scale spiking neural network is that updating all weights is quite expensive. Furthermore, after training, all information related to the computational task is hidden into the weight matrix, prohibiting us from a transparent understanding of circuit mechanisms. Therefore, in this work, we address these challenges by proposing a spiking mode-based training protocol, where the recurrent weight matrix is explained as a Hopfield-like multiplication of three matrices: input modes, output modes and a score matrix. The first advantage is that the weight is interpreted by input and output modes and their associated scores characterizing the importance of each decomposition term. The number of modes is thus adjustable, allowing more degrees of freedom for modeling the experimental data. This significantly reduces the training cost because of significantly reduced space complexity for learning. Training spiking networks is thus carried out in the mode-score space. The second advantage is that one can project the high dimensional neural activity (filtered spike train) in the state space onto the mode space which is typically of a low dimension, e.g., a few modes are sufficient to capture the shape of the underlying neural manifolds. We successfully apply our framework in two computational tasks—digit classification and selective sensory integration tasks. Our method thus accelerates the training of spiking neural networks by a Hopfield-like decomposition, and moreover this training leads to low-dimensional attractor structures of high-dimensional neural dynamics.
@@ I Introduction
We review relevant existing methods of training spiking neural networks and highlight the difference from ours. Recurrent neural networks (RNNs) using continuous rate dynamics of neurons can be used to generate coherent output sequences [7], in which the network weights are trained by the first-order reduced and controlled error (FORCE) method, where the inverse of the rate correlation matrix is iteratively estimated to update the weights during learning [7, 5]. This method can be generalized to supervised learning in spiking neural networks [8, 9, 10]. A recent work proposed to train the rate network first and then rescale the synaptic weights to adapt to a spiking setting [11], while earlier works proposed to map a trained continuous-variable rate RNN to a spiking RNN model [12, 13]. These methods rely either on standard efficient methods like FORCE for rate models, or on heuristic strategies to modify the weights on the rate counterpart. Another recently proposed route is a non-linear voltage-based three-factor learning rule for multilayer networks of deterministic integrate-and-fire neurons [14]. All these studies involve in real valued weights. In contrast, our current mode decomposition framework works directly on the mode space of the recurrent connections underlying the spiking dynamics, for which many biological plausible factors can be incorporated, e.g., membrane and synapse time scales, cell types, refractory time etc.
@@
In essence, we decompose the traditional real-valued weights as three matrices: the left one acts as the input mode space, the right one acts as the output mode space, and the middle one encodes the importance of each mode in the corresponding space. Therefore, we can interpret the weight as two mode spaces and a score matrix. The neural dynamics can thus be projected to the input mode space which is an intrinsically low dimensional space. In addition, the leaky integrated fire (LIF) model can be discretized into a discrete dynamics with intrinsic synaptic dynamics and threshold-type firing, for which a surrogate gradient descent with an additional steepness hyperparameter can be applied [6, 27]. We then show in experiments how the SMNN framework is successfully applied to classify pixel-by-pixel MNIST datasets and furthermore to context dependent computation in neuroscience. For both computational tasks, we reveal attractor structures in a low-dimensional mode space (in the number of modes) onto which the high dimensional neural dynamics can be projected. Overall, we construct a computational efficient (less model parameters required because of a few dominating modes) and conceptually simple framework to understand challenging spike-based computations (attractor picture in physics). This finding will inspire more physics stuides on the high dimensional spiking neural dynamics including those driven by learning.
@@ II.1 Recurrent spiking dynamics with mode-based connections
Before specifying the neural dynamics, we remark a few differences from the singular value decomposition method. First, the mode decomposition learning rule is inspired in physics by Hopfield model (a toy model of associative memory [30], see also our two recent papers [24, 26]). In the generalized form [26], one can flexibly interpret the rank number as the number of patterns, and this number is adjustable. In biology, this decomposition may implement a linear-nonlinear layered computation akin to dendritic integration [26]. Second, the elements of the importance matrix is not necessarily positive, and can be learned during the training protocol. Third, the input and output mode vectors are not necessarily orthogonal, allowing for more degrees of freedom to interpret experimental data or execute computational tasks. In addition, these mode vectors help to project the dynamics trajectories in the high-dimensional state space onto the low-dimensional mode space (the dimensionality is determined by $P\ll N$).
@@
where $\tau_{r}$ and $\tau_{d}$ refer to the synaptic rise time and the synaptic decay time, respectively; $t_{i}^{k}$ refers to the $k$-th spiking time of unit $i$. Different time scales of synaptic filters are due to different types of receptors (such as fast AMPA, relatively fast GABA, and slow NMDA receptors) with different temporal characteristics [1]. Note that the above synaptic filter has also been used in a previous work [11], which proposed to train the rate network first and then rescale the synaptic weights to adapt to the spiking setting. We remark that other types of synaptic filters can also be used, e.g. the simple exponential type used in previous works [9, 27].
@@ II.2.1 MNIST data learning
The MNIST dataset consists of handwritten digits, each of which is composed of 28x28 pixels, belonging to $10$ different classes (6000 images for each class). The dataset is commonly used as a benchmark classification for neural networks [32]. The handwritten digit is a static image, and thus must be transformed into spiking time series. We thus introduce an additional transformation layer consisting of spiking neurons before the hidden layer. More precisely, each image is converted into spiking activity using one neuron per pixel. The transformation layer is then modeled by a population of Poisson spiking neurons with a maximum firing rate $f_{\rm max}$. Each neuron in this layer generates a Poisson spike train with a rate $f_{i}=\frac{g_{i}}{255}f_{\rm max}$, where $g_{i}$ is the corresponding pixel intensity. We present the same static image every $0.2\,{\rm{ms}}$ (step size) and for a total of $100$ time steps as an input to the transformation layer (therefore for a duration $20\ {\rm{ms}}$), and this input image is converted into spiking activity as described above to the hidden layer [Figure 1 (a) and (b)].
@@ III Simulation results of mode decomposition learning
where a steepness parameter $k$ = 25. This is because, $\lim_{k\to\infty}\Theta^{\prime}(x)=\delta(x)$, where $x=U-U_{\rm thr}$. In the forward pass, the original form of the step function is used to model the spiking process. Other parameters include $\Delta t=0.2\,{\rm{ms}}$, $\tau_{\rm mem}=20\,{\rm{ms}}$, $\tau_{d}=30\,{\rm{ms}}$, $\tau_{r}=2\,{\rm{ms}}$, and the refractory period $t_{\rm ref}=2\,{\rm{ms}}$ for all tasks. These hyper-parameters are summarized in Table 3 (see Appendix C). Because we have not imposed sparsity constraint on the connectivity, we take the initialization scheme that $[\bm{\xi}^{\rm in}\bm{\Sigma}(\bm{\xi}^{\rm out})^{\top}]_{ij}\sim O(\frac{1}{\sqrt{N}})$, similar to what is done in multi-layered perceptron learning [26]. The training is implemented by the adaptive-moment-estimation (Adam [34]) based stochastic gradient descent algorithm that minimizes the loss function (cross-entropy for classification or mean-squared error for regression in the second task) with a learning rate of $0.001$. We remark that training in the mode space can be done by using automated differentiation on PyTorch. However, we still leave the detailed derivation of the learning rule to Appendix B. Codes are available in our GitHub [35].
@@ III.1 MNIST classification task
Figure 2: Learning performance of MNIST classification task. Five independent runs are used to estimate the standard deviation. (a) Test accuracy versus different mode size. Spiking model without mode decomposition learning (this counterpart is called SNN) is compared. SMNN indicates the mode-based learning of spiking networks. (b) Loss function as a function of training mini-batch. Each mini-batch is composed of $100$ digit images. The mode size varies, and the network size $N=100$. (c) Comparison of the accuracies between rate and spiking models with the same fixed network size $N$=200. Rate networks with mode-decomposition learning (MDL RNN) and without MDL (RNN) are also considered. (d) Membrane potential traces for three typical reservoir neurons in response to spike train inputs after training. The triangle or square marks when the filtered spike train takes a maximum value. $(P,N)=(3,100)$. Note that the displayed three moments for the maximal firing rates overlap with each other, but this is not the case for other neurons. (e) Spike trains of reservoir neurons with neuron firing rate (right) and population firing fraction (bottom) in response to an input image of digit $2$. $(P,N)=(3,100)$.
@@ III.2 Contextual integration task
As shown in Figure 3 (a), the test MSE [i.e., $\mathcal{L}^{2}$, see Eq. (11)] for the SMNN decreases with the mode size. In particular, a larger neural population is better. Even for a few modes like $P=3$, the predicted output already matches the target very well [see Figure 3 (c)]. The SNN without the mode decomposition achieves a lower MSE, but we emphasize that the accuracies behind the MSE within the shown range (from $0.01$ to $0.02$) do not look very different [see Figure 3 (c) for the case of $P=3$]. In this sense, compared to SNN, SMNN saves the training parameters. The training dynamics in Figure 3 (b) shows that a larger value of $P$ speeds up the learning. Figure 3 (c) confirms that our training protocol succeeds in reproducing the result of Monkey’s experiments on the task of flexible selective integration of sensory inputs. We can even look at the dynamics profile of the membrane potential [Figure 3 (d)]. During the stimulus period, each neuron has its own time scale to encode the input signals, and at a network scale, we do not observe any regular patterns but an asynchronous pattern typically observed in prefrontal cortex. After the sensory input is turned off, the network is immediately required to make a decision, and we observe that the firing frequency of the neural pool is elevated already a bit earlier than the moment of removing input, but a very low population firing rate is still maintained [Figure 3 (e)]. This observation is a reflection of neural dynamics on the neural manifolds underlying the perceptual decision making, which we shall detail below.
@@ III.3 Power law for connectivity importance scores
Figure 4: Connectivity importance $|\lambda_{\mu}|$ versus rank (in descending order). Both MNIST classification and contextual integration task are considered, and the simulation conditions are the same with that in Figure 2 for MNIST and that in Figure 3 for context dependent computation. We fix $P=30$. Three independent runs are used to estimate the standard deviation. We also define a more precise measure $\tau_{\mu}=\chi\|\bm{\xi}^{\rm in}_{\mu}\|_{2}+\chi\|\bm{\xi}^{\rm out}_{\mu}\|_{2}+|\lambda_{\mu}|$, where $\chi=\sum_{\mu}|\lambda_{\mu}|/\sum_{\mu}(\|\bm{\xi}^{\rm in}_{\mu}\|_{2}+\|\bm{\xi}^{\rm out}_{\mu}\|_{2})$. (a,b) MNIST classification. (c,d) Contextual integration task. There appears the piecewise power law behavior for the $\tau$-measure in the log-log plot (b,d). The colored vertical dashed lines mark the fitting ranges for different network sizes.
@@ III.4 Reduced dynamics in the mode space
Here, we will show how the dynamics of network activity based on our SMNN learning can be projected to a low-dimensional counterpart. We follow the previous works of low-rank recurrent neural networks [21, 22]. To construct a subspace spanned by orthogonal bases, we first split each column of $\bm{W}^{\rm in}$ ($\in\mathbb{R}^{N\times N_{\rm in}}$) into the parts parallel and orthogonal to the input mode $\bm{\xi}^{\rm in}$ (or the output mode),
@@ III.6 Spiking variability measured by Fano factor
The spiking activity is commonly irregular, displaying strong variability even when the neural population is shown identical stimuli. One can thus count the spiking events in a time window of duration $T$ for one neuron, and then repeat the experiment over many trials and finally measure the mean and variance of the counts. The spiking variability is measured by the Fano factor calculated as [1]
@@
We plot the Fano factor distribution across neurons in Fig. 7. The Fano factor is estimated over the entire experimental window. For the MNIST classification task, the Fano factor is mainly distributed from $2$ to $4$ [Figure 7 (a)], indicating highly unreliable firing patterns, which corresponds to that observed in Figure 2 (e). For an input image, some neurons fire at high frequencies, while others remain silent. This non-Poissonian phenomenon has also been observed in cortical areas [36]. For the context-integration task, the Fano factor is surprisingly concentrated around one [Figure 7 (b)], except for a long tail at large values of Fano factor. This model observation highlights the potential of our method in studying the alignment between the simulation and neural experiments of context dependent cognitive decision making.
@@ III.7 Ablation studies on the refractory, activation function and network robustness
where we set $\beta=20$ to mimic the step function. Note that in this control experiment, we use this transfer function both in forward and backward passes during the training. As shown in Figure 8 (a) for the MNIST task, the absence of the refractory period allows the network to respond more correctly to external stimuli. Performance without the refractory period is slightly better than with the refractory period, but as the mode number increases, the gap shrinks. For the contextual integration task, an appropriate refractory period is able to optimize network performance [Figure 8 (b)], consistent with previous results in standard SNNs [11]. Finally, we remark that networks that use a smooth transfer function instead of the step function perform much worse than spiking networks using a sharp step function (neurons can thus fire or be silent), indicating the importance of spiking event.
@@
We next consider the network robustness in a weight-pruning experiment. More precisely, for a well-trained connectivity matrix $W^{\rm rec}_{ij}=\sum_{\mu}\lambda_{\mu}\xi^{\rm in}_{i\mu}\xi^{\rm out}_{j\mu}$, we set some elements where their absolute values $|W^{\rm rec}_{ij}|<\theta$ to zero, where $\theta$ is an adjustable pruning threshold. This operation leads to a diluted connectivity matrix $\hat{\bm{W}}^{\rm rec}$:
@@
Finally, we also carry out the pruning experiment on trained SNNs where the mode decomposition is not used. As shown in Figure 10, the SMNN is able to maintain a better accuracy with fewer parameters compared to a well-trained SNN subject to an equivalent pruning after training.
Figure 10: Test performance in pruning experiments on trained SNNs ($N=100$) where the weights are set to zero if its absolute value falls below a threshold. (a) Test accuracy for the MNIST classification task. (b) Test mean square error for the contextual integration task. The vertical line marks the model complexity of SMNN [$(1-(2NP+P)/N^{2})$×$\times 100\%$], and the corresponding test performances are specified in the brackets.